OpenStudy (anonymous):

The Steamboat aloftser in Yellowstone National Park, Wyoming is capable of shooting its hot water up from the ground with a speed of 48.0 m/s. How high can this aloftser shoot? (HINT: USE KINEMATIC EQUATIONS)

OpenStudy (anonymous):

1/2 m v^2= m g h 1/2 v^2 = g h 1/2 v^2 ------- =h g

OpenStudy (anonymous):

Um... could you explain please because it doesnt look like something I learned before

OpenStudy (anonymous):

There is something called conservation of energy. So in the beginning all the energy were kinetic(moving energy) = 1/2 m v^2 when it is at the top , all the energy is potential() = mgh they should be the same because of energy conservation 1/2 mv^2 = m gh mass(m) cancels out

OpenStudy (anonymous):

1/2 v^2= gh 1/2 v^2 ------- =h g

OpenStudy (anonymous):

Or you can use one of the kinetic equaton \[(v_f)^2=(v_i)^2 - 2 g h\] at the top \[v_f=0\] \[0=(v_i)^2 - 2 g h\] \[h=(v_i)^2/2g \]